OpenStudy (anonymous):

fan and medal Solve for x: −3|2x + 6| = −12 x = 1 and x = 5 x = −1 and x = −5 x = −9 and x = 3 No solution

OpenStudy (compassionate):

Hi, [(-5, 0), (-1, 0)]

OpenStudy (anonymous):

@Compassionate , thank you.

OpenStudy (the_fizicx99):

−3|2x + 6| = −12 Divide both sides by -3 |2x + 6| = 4 Subtract 6 from both sides, 2x = -2 /2 /2 x = -1 Now make a new equation except with the answer being negative, |2x + 6| = -4 -6 -6 2x = -10 /2 /2 x = -5 (-1,0) & (-5,0)